# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 4.3 Expansion of the asymptotic BAR
$\quad=(1-e^{\theta_{1}\ell_{0}})\frac{r}{\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+r2e^{\theta_{1}\ell_{0}}\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r),$
@@
$\quad=\frac{re^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})}{(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r)$
$\quad=\frac{rb_{1}\mu\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\big{(}e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+r\frac{\beta_{1}e^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})+\theta_{1}e^{\theta_{1}\ell_{0}}(e^{\beta_{1}\ell_{0}}-1)}{(e^{\beta_{1}\ell_{0}}-1)}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r).$
@@
$\psi^{(r)}_{1}(\theta_{1})$ $=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}-\theta_{1}e^{\theta_{1}\ell_{0}}+(\theta_{1}-\beta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
$=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}$
$\quad+\frac{\beta_{1}(e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}-(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2e^{\theta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
@@
$\psi^{(r)}_{2}(\theta_{2})=\frac{e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\beta_{2}\varphi^{(r)}_{2}(0)+(\beta_{2}-(\beta_{2}-\theta_{2}))\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}+\theta_{2}o(r),$
@@
then the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{1}(\theta_{2})$ can be obtained as long as the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{1}(0)$ exist. This suggests that (5.1) must be a key condition for Theorem 3.1. Taking this into account, we separate the proof of Theorem 3.1 in two steps by two propositions.
@@
In Theorem 3.1, the assumption (3.f) can be replaced by (5.1).
@@
Under the assumptions (3.a)–(3.e), (5.1) holds if
$\lim_{r\downarrow 0}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})=0.$ (5.2)
Furthermore, (5.2) holds if (3.f) holds. Hence, (3.f) implies (5.1).
In the next two subsections, we prove these propositions. For their proofs, we will use the following lemmas, which are proved in Appendix A.5, respectively.
@@
Under the assumptions (3.a)–(3.e), if
$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,$ (5.3)
then
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.4)
@@
Under the assumptions (3.a)–(3.e), if (5.4) holds, then
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.5)
@@
From the proof of Proposition 5.2, we can see that (5.4) and (5.5) imply (5.2), and therefore (5.3) is sufficient for (5.2) and (5.1). We think that (5.3) is likely true because $L^{(r)}(\cdot)$ would weakly converge to a reflected diffusion and the stationary distribution of the reflected diffusion has a density. However, the verification of (5.3) would need a process limit for $L^{(r)}(\cdot)$, so this verification is beyond the present framework.
@@
Throughout this proof, the part (i), (ii) and (iii) of Theorem 3.1 under the condition (5.1) instead of (3.f) are denoted by the part (i)’, (ii)’ and (iii)’, respectively.
We first prove the part (i)’. Since the condition in the part (i)’ is given by (3.6), Corollary 4.3 implies that
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (5.7)
$\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{0}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (5.8)
@@
We next prove the part (ii)’. For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumptions in the part (ii)’. First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
@@
$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (5.9)
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}b_{1}\mu\varphi^{(r)}_{1}(0)+o(r),&b\not=0,\\ \end{cases}$ (5.10)
@@
$\lambda_{2}\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
@@
$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$ (5.12)
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}\sigma_{1}^{2}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.13)
@@
$\frac{\lambda_{2}b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$ (5.14)
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.15)
@@
$\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(\theta_{1})=\widetilde{\varphi}_{1}(\theta_{1})\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(\theta_{1})\overline{\varphi}_{1}(0).$ (5.16)
Since the distribution $\widetilde{\nu}_{1}$ has the bounded density $h_{1|cd}$ on $[0,\ell_{0}]$, this implies that
$\lim_{n\to\infty}\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]\leq\lim_{n\to\infty}\mathbb{P}[\overline{r}_{n}L^{(\overline{r}_{n})}\in[\ell_{0}-\varepsilon,\overline{r}_{n}\ell^{(\overline{r}_{n})}_{0}]]$
@@
$\alpha^{(\overline{r}_{n})}_{e}(\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2})=\frac{rc_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)-r\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)+o(\overline{r}_{n}),$
@@
$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)=\mu b_{2}(1-\varphi^{(\overline{r}_{n})}_{1}(0))+o(1).$
@@
$\frac{b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(\overline{r}_{n})}_{1}(0)-\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)=o(1).$
@@
$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
@@
Using (5.2), we prove (5.1) by reductio ad absurdum. First, assume that ${\mathcal{E}}^{(r)}_{1}$ does not vanish as $r\downarrow 0$. Then, there are a sequence $\{r_{n}\in(0,1];n\geq 0\}$ and a constant $\delta_{1}\not=0$ such that $r_{n}\downarrow 0$ and ${\mathcal{E}}^{(r_{n})}_{1}\to\delta_{1}$ as $n\to\infty$, because ${\mathcal{E}}^{(r)}_{1}$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
@@
$\lim_{n\to\infty}\varphi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}\not=0.\end{cases}$
This shows that $\nu^{(r)}_{1}$ weakly converges to a finite measure on $[0,\ell_{0}]$, but this measure cannot have a negative mass at $\ell_{0}$, so we must have that $\delta_{1}<0$ because $\delta_{1}\not=0$.

### v8

@@ 4.3 Expansion of the asymptotic BAR
$\quad=(1-e^{\theta_{1}\ell_{1}})\frac{r}{\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+r2e^{\theta_{1}\ell_{1}}\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r),$
@@
$\quad=\frac{re^{\beta_{1}\ell_{1}}(1-e^{\theta_{1}\ell_{1}})}{(e^{\beta_{1}\ell_{1}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+re^{\theta_{1}\ell_{1}}\big{(}b_{1}\mu\varphi^{(r)}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r)$
$\quad=\frac{rb_{1}\mu\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{1}}-1}\big{(}e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}})+r\frac{\beta_{1}e^{\beta_{1}\ell_{1}}(1-e^{\theta_{1}\ell_{1}})+\theta_{1}e^{\theta_{1}\ell_{1}}(e^{\beta_{1}\ell_{1}}-1)}{(e^{\beta_{1}\ell_{1}}-1)}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r).$
@@
$\psi^{(r)}_{1}(\theta_{1})$ $=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}$
$\quad+\frac{\beta_{1}e^{\beta_{1}\ell_{1}}-\theta_{1}e^{\theta_{1}\ell_{1}}+(\theta_{1}-\beta_{1})e^{\theta_{1}\ell_{1}}e^{\beta_{1}\ell_{1}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
$=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}$
$\quad+\frac{\beta_{1}(e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}})+(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{1}}-(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{1}}e^{\beta_{1}\ell_{1}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2e^{\theta_{1}\ell_{1}}}{c_{1}\mu\sigma_{1}^{2}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
@@
$\psi^{(r)}_{2}(\theta_{2})=\frac{e^{\theta_{2}\ell_{1}}}{\beta_{2}-\theta_{2}}\Big{(}\beta_{2}\varphi^{(r)}_{2}(0)+(\beta_{2}-(\beta_{2}-\theta_{2}))\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}+\theta_{2}o(r),$
@@
then the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{1}(\theta_{2})$ can be obtained as long as the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{1}(0)$ exist. This suggests that (5.1) must be a key condition for Theorem 3.1. Taking this into account, we separate the proof of Theorem 3.1 in two steps.
@@
If (5.1) holds, then Theorem 3.1 holds without the assumption (3.f).
@@
Under the assumptions (3.a)–(3.e), if
$\lim_{r\downarrow 0}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})=0,$ (5.2)
then (5.1) holds. In particular, (3.f) implies (5.2) and therefore (5.1).
These propositions are proved in the next two subsections. Obviously, they prove Theorem 3.1. To prove Proposition 5.2, we will use the following lemma, which is proved in Appendix A.6.
@@
Under the assumptions (3.a)–(3.e), we have
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{i}1(L^{(r)}(0)=\ell^{(r)}_{1})]=0,\qquad i=1,2,$ (5.3)
if either $\lambda_{2}\not=\lambda_{2}$ or
$\lim_{r\downarrow 0}\mathbb{P}^{(r)}[L^{(r)}=\ell^{(r)}_{1}]=0.$ (5.4)
@@
Under the assumptions (3.a)–(3.e), (5.3) implies (5.2).
@@
By Lemmas 5.1 and 5.2 and Proposition 5.2, (5.4) is sufficient for (5.1). We conjecture that (5.4) is true because $L^{(r)}(\cdot)$ would weakly converge to a reflected diffusion whose stationary distribution has a density. However, the verification of (5.4) would need a process limit for $L^{(r)}(\cdot)$, so it is beyond the present framework.
@@
Throughout this proof, the parts (i), (ii) and (iii) of Theorem 3.1 under the condition (5.1) instead of (3.f) are denoted by the part (i)’, (ii)’ and (iii)’, respectively.
We first prove the part (i)’. Since (3.6) holds, Corollary 4.3 implies that
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{1}}-1}{\ell_{1}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (5.6)
$\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{1}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (5.7)
@@
We next prove the part (ii)’. For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumptions in the part (ii)’. First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$ by Lemma 2.1 and Theorem 3.1 of [12], so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)\;=\ell^{(r)}_{1})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]+o(r).$
@@
$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{1})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}])=o(r).$ (5.8)
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{1}}c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{1}}-1}b_{1}\mu\varphi^{(r)}_{1}(0)+o(r),&b\not=0,\\ \end{cases}$ (5.9)
@@
$\lambda_{2}\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
@@
$\frac{\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$ (5.11)
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{2b_{2}\ell_{1}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{1}}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}\sigma_{1}^{2}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{1}}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.12)
@@
$\frac{\lambda_{2}b_{1}\mu}{e^{\beta_{1}\ell_{1}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$ (5.13)
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.14)
@@
$\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(\theta_{1})=\lim_{n\to\infty}\psi^{(\overline{r}_{n})}_{1}(\theta_{1})=\widetilde{\varphi}_{1}(\theta_{1})\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(\theta_{1})\overline{\varphi}_{1}(0).$ (5.15)
Since the distribution $\widetilde{\nu}_{1}$ has the bounded density $h_{1|cd}$ on $[0,\ell_{1}]$, this implies that
$\lim_{n\to\infty}\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]\leq\lim_{n\to\infty}\mathbb{P}\left[\overline{r}_{n}L^{(\overline{r}_{n})}\in[\ell_{1}-\varepsilon,\overline{r}_{n}\ell^{(\overline{r}_{n})}_{0}]\right]$
@@
$\alpha^{(\overline{r}_{n})}_{e}(\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2})=\frac{rc_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(\overline{r}_{n})}_{1}(0)-r\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)+o(\overline{r}_{n}),$
@@
$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(\overline{r}_{n})}_{1}(0)=\mu b_{2}(1-\varphi^{(\overline{r}_{n})}_{1}(0))+o(1).$
@@
$\frac{b_{1}\mu}{e^{\beta_{1}\ell_{1}}-1}\varphi^{(\overline{r}_{n})}_{1}(0)-\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)=o(1).$
@@
$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
@@
We prove that (5.2) implies (5.1) by reductio ad absurdum. First, assume that ${\mathcal{E}}^{(r)}_{1}$ does not vanish as $r\downarrow 0$. Then, there are a sequence $\{r_{n}\in(0,1];n\geq 0\}$ and a constant $\delta_{1}\not=0$ such that $r_{n}\downarrow 0$ and ${\mathcal{E}}^{(r_{n})}_{1}\to\delta_{1}$ as $n\to\infty$, because ${\mathcal{E}}^{(r)}_{1}$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
@@
$\lim_{n\to\infty}\varphi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{1}}-1}{\ell_{1}\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}},&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}},&b_{1}\not=0.\end{cases}$
This shows that $\nu^{(r)}_{1}$ weakly converges to a finite measure on $[0,\ell_{1}]$, but this measure cannot have a negative mass at $\ell_{1}$, so we must have that $\delta_{1}<0$ because $\delta_{1}\not=0$.